OpenStudy (anonymous):

Find all points of extrema on the interval [0,2pi] if y=x-cosx

OpenStudy (anonymous):

So first differentiaite it y'=1+sin x to find all critical points put y'=0 you get sinx=-1 which is true at x=pi only in the given interval to check whether it is maximum you should ahve y''<0 but y''=cosx at x=pi cosx=0 hence y''=0 which does not satisfy the condtion y''<0 Hence there is no extreme point in the given interval....

OpenStudy (anonymous):

You do know differentiation and its applications right?

OpenStudy (turingtest):

what is the derivative of x-cosx ?

OpenStudy (anonymous):

yes, yes i do

OpenStudy (turingtest):

what is \[\frac{d}{dx}(x-\cos x)\] ?

OpenStudy (anonymous):

sinx?

OpenStudy (anonymous):

nvm its 1-sinx

OpenStudy (turingtest):

not quite remember derivatives of co-functions (cosine, cotangent, csc) are negative\[\frac{d}{dx}(x-\cos x)=\frac{d}{dx}x-\frac{d}{dx}\cos x=1+\sin x\]

OpenStudy (anonymous):

oh ok so i equal that to zero then?

OpenStudy (turingtest):

right what do you get for x?

OpenStudy (anonymous):

-pi over 2

OpenStudy (turingtest):

but that's not in the interval [0,2pi] now is it?

OpenStudy (turingtest):

what is the equivalent angle to -(pi/2) that is in the interval [0,2pi] ?

OpenStudy (anonymous):

this is the unit circle correct?

OpenStudy (turingtest):

yes go around it the other way and get to the same point as -pi/2 what do you call that point going the other way (counterclockwise)?

OpenStudy (anonymous):

0,1?

OpenStudy (turingtest):



OpenStudy (turingtest):

clockwise we get